# TC-GS: A Faster Gaussian Splatting Module Utilizing Tensor Cores

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) renders pixels by rasterizing Gaussian primitives, where conditional alpha-blending dominates the time cost in the rendering pipeline. This paper proposes TC-GS, an algorithm-independent universal module that expands Tensor Core (TCU) applicability for 3DGS, leading to substantial speedups and seamless integration into existing 3DGS optimization frameworks. The key innovation lies in mapping alpha computation to matrix multiplication, fully utilizing otherwise idle TCUs in existing 3DGS implementations. TC-GS provides plug-and-play acceleration for existing top-tier acceleration algorithms tightly coupled with rendering pipeline designs, like Gaussian compression and redundancy elimination algorithms. Additionally, we introduce a global-to-local coordinate transformation to mitigate rounding errors from quadratic terms of pixel coordinates caused by Tensor Core half-precision computation. Extensive experiments demonstrate that our method maintains rendering quality while providing an additional 2.18× speedup over existing Gaussian acceleration algorithms, thus reaching up to a total 5.6× acceleration. The code are currently available at anonymous https://github.com/TensorCore3DGS/3DGSTensorCore.
@@
Figure 1: TC-GS: a Tensor Core-based acceleration module for 3D Gaussian Splatting that can be effortlessly applied to existing 3DGS rendering pipelines. Integrated into 3DGS and its variants, it achieves an extra 2x speedup, resulting in an overall 4× acceleration while preserving rendering quality.
Keywords Gaussian Splatting  $\cdot$ Tensor Cores  $\cdot$ Real-Time Rendering
1 Introduction
@@
3D Gaussian Splatting (3DGS) represents a significant advancement in neural rendering, employing Gaussian primitives to achieve NeRF-comparable quality while reducing optimization time to tens of minutes [1, 2]. However, 3DGS rendering speed is heavily constrained by heavy model parameters and inefficient rendering pipelines[3, 4, 5]. Further acceleration remains essential for deployment on resource-constrained devices, real-time large-scale scene rendering systems, and low-latency edge computing scenarios.
@@
Despite numerous efforts to enhance computational efficiency [6, 7, 8, 9], they fail to tap into the most powerful computing units in modern GPUs. Moreover, existing acceleration modules are tightly coupled with core components, preventing modular reuse and cross-method adaptation. This coupling further limits the practical deployment and scalability of 3DGS-based solutions.
@@
Recent research [6, 7, 8, 9] has identified the conditional alpha-blending process as the primary pipeline bottleneck, which composites pixel colors with depth-sorted splatted Gaussian primitives. This process comprises three steps: (1) Alpha computation: computing the alpha value (opacity) of each splatted Gaussian fragment on the pixel; (2) Culling: eliminating fragments with low alpha values; and (3) Blending: compositing remaining fragments into final pixel colors. Many approaches [10, 6, 11] employ early culling strategies to reduce the range of splatted Gaussians, thereby minimizing redundant alpha-blending computations. Although eliminating redundant tile computation, they require corresponding pipeline designs to mitigate load imbalance issues introduced by precise intersection [7], or sacrifice certain accuracy [6]. Additionally, we observe that over 80% of Gaussian-pixel pairs requiring culling still remain in the pipeline. Consequently, the first two steps are the dominant part of the alpha-blending process.
@@
Unlike previous efforts to remove redundant computing [6, 7, 8, 9], we focus on utilizing Tensor Core Units (TCUs) for accelerating alpha computation, which remains an unexplored area. In modern GPUs, TCUs are deployed to perform matrix multiply accumulate (MMA), i.e., $D=A\times B+C$ operation within a single clock cycle, achieving high throughput on general matrix multiplication (GEMM) based computations including MLPs, CNNs, and transformers [12]. However, due to their specialized computation pattern, TCUs struggle to be applied to non-GEMM computations, leaving them idle in 3DGS.
@@
To tackle this problem, this paper proposes TC-GS, a hardware-aligned redesign of alpha computation that unlocks Tensor Core acceleration for all 3DGS algorithms. TC-GS formulates a mapping from pixels and Gaussians into two matrices, where their matrix product yields logarithmic alpha values in batches. Implemented as a standalone module, TC-GS not only significantly improves rendering efficiency but also enables seamless integration with existing 3DGS acceleration kernels, further enhancing overall performance.
@@
  • Our detailed runtime analysis of the rendering pipeline proves that alpha computation in conditional alpha-blending is the performance bottleneck of the 3DGS inference rendering.
  • TC-GS is the first to expand Tensor Core applicability for 3DGS, through mapping the alpha computation into matrix multiplication, enabling full utilization of otherwise idle TCUs in existing 3DGS implementations. As an algorithm-independent acceleration module, TC-GS seamlessly integrates into diverse 3DGS acceleration kernels and frameworks to achieve more efficient rendering performance.
@@
  • We introduce a global-to-local coordinate transformation (G2L) that reduces the absolute magnitudes of quadratic coordinate terms, which are sensitive to Tensor Core half-precision computation, enabling lossless rendering quality while maintaining full acceleration benefits.
2 Related Works
2.1 Fast 3D Gaussian Splatting
@@
3DGS [13] has emerged as a leading real-time rendering technique with its photo-realistic visual quality, finding a wide range of applications including autonomous driving [14], robotics [15], virtual reality [5]. Current acceleration research explores two primary optimization strategies to enhance 3DGS performance. The first strategy involves modifying the algorithm itself to achieve better efficiency [10, 1, 16]. CompactGS [16] and gsplat [10] introduce more efficient radiance field representations and adaptive resolution primitive pruning algorithms, while Taming-3DGS [1] presents mathematically equivalent but computationally efficient solutions for gradient computation and attribute updates, substantially accelerating training speed. The second category focuses on improving computational efficiency of the original algorithm through intelligent task scheduling and CUDA kernel optimization [7, 8, 6, 10, 17, 2]. Balanced-3DGS [17] optimizes the forward computation of rendering CUDA kernels within warps while minimizing thread idle time and maximizing resource utilization by evenly distributing tasks across compute blocks. Several works, including AdRGaussian [6], FlashGS [7], and SpeedySplat [8], aim to reduce the number of pixels processed per Gaussian by designing more precise Gaussian-tile bounding boxes, with FlashGS [7] and SpeedySplat [8] further computing exact Gaussian-tile intersections. DISTWAR [2] enhances processing efficiency through sophisticated thread management strategies, fully leveraging warp-level reductions in SM sub-cores and intelligently distributing atomic computations between SM and L2 atomic units based on memory access patterns. While these methods significantly accelerate training or rendering speed, strong coupling between algorithmic modules hinders transferability and reusability. Moreover, they fail to leverage modern hardware acceleration features such as Tensor Cores, further limiting the practical deployment and scalability of 3DGS solutions.
2.2 Tensor Cores for Non-GEMM Based Computations
Tensor Cores Units are specialized hardware units developed by NVIDIA to accelerate Deep Learning applications [18]. Each Tensor Core can perform a matrix multiply accumulate operation ${D=A\times B+C}$ on small matrices within a GPU cycle, where A and B must be in half percision format while the accumulators, C and D, can be either single or half percision. Tensor Core can achieve higher throughput than normal CUDA Cores: for instance on the NVIDIA A100 GPU, CUDA Cores has maximum 78 TFLOPS fp16 performance, whereas Tensor Cores can achieve 312 TFLOPS which is 4$\times$ times faster than CUDA Cores.
@@ 2.2 Tensor Cores for Non-GEMM Based Computations
Despite this theoretical advantage, it is not a trival task to exploit Tensor Cores in arbitrary applications since Tensor Cores only support the matrix–multiply–accumulate (MMA) instruction rather than the full range of CUDA instructions set. This restriction demands that algorithms be reformulated in terms of small matrix operations, but the reward is exceptionally fast execution. Consequently, the strict requirements yet remarkable speed of Tensor Core operations have attracted researchers to explore adapting non-machine learning algorithms to GPU Tensor Cores, leveraging their performance potential. Previous work has efficiently mapped reduction and scan primitives to Tensor Cores[19, 20].Recently ConvStencil[21] has managed to adapt stencil computation to TCU. However, to our best knowledge, no prior work has explored how to exploit Tensor Cores directly within a 3D Gaussian Splatting rendering pipeline.
@@
Table 1: Tensor Core & CUDA Core Comparision among various GPUs. * stands for using new sparse feature
@@
H100 (PCIe) 96 TFLOPS 800 TFLOPS | 1600 TFLOPS*
A100 78 TFLOPS 312 TFLOPS | 624 TFLOPS*
@@
RTX 4090 82.6 TFLOPS 330.3 TFLOPS | 660.6 TFLOPS*
3 Preliminaries and Observations
3.1 3DGS Rendering Pipeline
A 3DGS model consists of a set of 3D Gaussians primitives $\mathcal{G}=\{G^{(1)},G^{(2)},\cdots,G^{(P)}\}$. Each Gaussian $G:\mathbb{R}^{3}\to\mathbb{R}$ is parameterized by its mean $\mu\in\mathbb{R}^{3}$, semi-definite covariance $\Sigma\in\mathbb{R}^{3\times 3}$, opacity $o\in(0,1]$, and color $c\in[0,1]^{3}$ and maps any 3D position $x\in\mathbb{R}^{3}$ to its density:
$G(x)=o\mathrm{e}^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)}.$ (1)
@@ 3 Preliminaries and Observations
The preprocess stage projects 3D Gaussians into 2D planes in parallel according to the viewing transform $W$ of a given camera:
$\mu^{\prime}=W\mu\in\mathbb{R}^{2},\quad\Sigma^{\prime}=JW\Sigma W^{T}J^{T}\in\mathbb{R}^{2\times 2},$ (2)
where $J=\frac{\partial{\mu^{\prime}}}{\partial\mu}$ is the Jacobian of projection $W$. The complexity of the preprocess stage is $\mathcal{E}_{p}=O(P)$, where $P$ is the number of Gaussians.
@@
For each projected Gaussian, this stage identifies all the tiles it covers. Then, for each tile $\mathcal{T}\in\mathcal{S}$, the overlapped Gaussians are sorted by depth. A tile $\mathcal{T}$ coupled with a covering Gaussian $G^{\prime}$ is denoted as a splat $(\mathcal{T},G^{\prime})$. The total number of splats given by
$N=\sum_{i=1}^{P}\sum_{\mathcal{T}\in\mathcal{S}}\mathbf{1}(\mathcal{T}\cap G^{\prime(i)}\neq\emptyset)=\sum_{\mathcal{T}\in\mathcal{S}}n(\mathcal{T}),$ (3)
where $\mathbf{1}$ is the indicator function that checks whether the Gaussian covers the tile. The process of determining coverage depends on the specific algorithm used. The complexity of the sorting stage is $\mathcal{E}_{s}=O(dN)$, where $d=64$ is the coefficient in radix-sort.
Input: Tile $\mathcal{T}$, Sorted Gaussians $\mathcal{G^{\prime}(\mathcal{T})}=(G^{\prime(1)},\cdots,G^{\prime(n(\mathcal{T}))})$
@@
for pixel $p\in\mathcal{T}$ in parallel do
@@
This stage renders the final color of each pixel in a tile $\mathcal{T}=(p_{1},p_{2},\cdots,p_{|\mathcal{T}|})$ in parallel as Algorithm 1. Initially, for each projected overlapping Gaussian, the renderer will compute the local opacity $\alpha_{i}^{(j)}$ at each pixel $p_{i}=(p_{x_{i}},p_{y_{i}})$:
$\alpha_{i}^{(j)}=o^{(j)}\exp{\left(-\frac{1}{2}(\mu^{\prime(j)}-p_{i})^{T}(\Sigma^{\prime(j)})^{-1}(\mu^{\prime(j)}-p_{i})\right)}.$ (4)
@@
$\alpha_{i}^{(j)}<\dfrac{1}{255},$ (5)
Otherwise, the surviving Gaussians will contribute the pixel’s color:
$C_{i}=\sum_{j=1}^{n(\mathcal{T})}c^{(j)}\alpha_{i}^{(j)}T_{i}^{(j)},\quad T_{i}^{(j)}=\prod_{k=1}^{j-1}(1-\alpha_{i}^{(k)}).$ (6)
where $T_{i}^{(j)}$ denotes the transmissivity and $n(\mathcal{T})$ is the number of splats of the tile $\mathcal{T}$. If $T_{i}^{(j)}<0.0001$, the renderer will terminate the blending process for that pixel and output the color. A pixel $p$ rendered by a Gaussian $G$ is referred to as a fragment, $(p,G)$.

### v2

@@ Abstract.
3D Gaussian Splatting (3DGS) renders pixels by rasterizing Gaussian primitives, where conditional alpha-blending dominates the computational cost in the rendering pipeline. This paper proposes TC-GS, an algorithm-independ-ent universal module that expands the applicability of Tensor Core (TCU) for 3DGS, leading to substantial speedups and seamless integration into existing 3DGS optimization frameworks. The key innovation lies in mapping alpha computation to matrix multiplication, fully utilizing otherwise idle TCUs in existing 3DGS implementations. TC-GS provides plug-and-play acceleration for existing top-tier acceleration algorithms and integrates seamlessly with rendering pipeline designs, such as Gaussian compression and redundancy elimination algorithms. Additionally, we introduce a global-to-local coordinate transformation to mitigate rounding errors from quadratic terms of pixel coordinates caused by Tensor Core half-precision computation. Extensive experiments demonstrate that our method maintains rendering quality while providing an additional 2.18× speedup over existing Gaussian acceleration algorithms, thereby achieving a total acceleration of up to 5.6×.
Gaussian Splatting, Tensor Cores, Real-Time Rendering
submissionid: 2424journalyear: 2025copyright: ccconference: arXiv Preprint Version; October 11th; 2025ccs: Computing methodologies Rasterizationccs: Computing methodologies Parallel algorithms
@@
Figure 1. TC-GS: a Tensor Core-based acceleration module for 3D Gaussian Splatting that can be effortlessly applied to existing 3DGS rendering pipelines. Integrated into 3DGS and its variants, it achieves an extra $2\times$ speedup while preserving rendering quality.:
1. Introduction
@@
3D Gaussian Splatting (3DGS) represents a significant advancement in neural rendering, employing Gaussian primitives to achieve NeRF-comparable quality while reducing optimization time to tens of minutes (Mallick et al., 2024; Durvasula et al., 2023). However, 3DGS rendering speed is heavily constrained by extensive model parameters and inefficient rendering pipelines(Lu et al., 2024; Ren et al., 2024; Jiang et al., 2024). Further acceleration remains essential for deployment on resource-constrained devices, real-time large-scale scene rendering systems, and low-latency edge computing scenarios.
@@
Despite numerous efforts to enhance computational efficiency (Wang et al., 2024; Feng et al., 2024; Hanson et al., 2024; Huang et al., 2025), they fail to tap into the most powerful computing units in modern GPUs. Moreover, existing acceleration modules are tightly coupled with core components, preventing modular reuse and inter-method adaptation. This coupling further limits the practical deployment and scalability of 3DGS-based solutions.
@@
Recent research (Wang et al., 2024; Feng et al., 2024; Hanson et al., 2024; Huang et al., 2025) has identified the conditional alpha-blending process as the primary pipeline bottleneck. This process, which composites pixel colors with depth-sorted splatted Gaussian primitives, comprises three steps: (1) Alpha computation: computing the alpha value (opacity) of each splatted Gaussian fragment on the pixel; (2) Culling: eliminating fragments with low alpha values; and (3) Blending: compositing remaining fragments into final pixel colors.
Many approaches (Ye et al., 2025; Wang et al., 2024; Fan et al., 2024) employ early culling strategies to reduce the range of splatted Gaussians, thus minimizing redundant alpha-blending computations. Although they eliminate redundant computation, these methods require corresponding pipeline designs to mitigate load imbalance issues introduced by precise intersection (Feng et al., 2024) calculations or sacrifice certain precision(Wang et al., 2024). Additionally, we observe that over 80% of Gaussian-pixel pairs that require culling remain in the pipeline. Consequently, the first two steps are the dominant part of the alpha-blending process.
@@
Unlike previous efforts to remove redundant computing (Wang et al., 2024; Feng et al., 2024; Hanson et al., 2024; Huang et al., 2025), we focus on utilizing Tensor Core Units (TCUs) to accelerate alpha computation, which remains an unexplored area. In modern GPUs, TCUs are deployed to perform matrix multiply accumulate (MMA), i.e., $D=A\times B+C$ operations within a clock cycle, achieving high throughput on general matrix multiplication (GEMM) based computations including MLPs, CNNs, and transformers (Vaswani et al., 2017). However, due to their specialized computation pattern, TCUs struggle to be applied to non-GEMM computations, leaving them idle in 3DGS.
@@
To tackle this problem, this paper proposes TC-GS, a hardware-aligned redesign of alpha computation that unlocks Tensor Core acceleration for all 3DGS algorithms. TC-GS formulates a mapping from pixels and Gaussians into two matrices, where their matrix product yields logarithmic alpha values in batches. Implemented as a standalone module, TC-GS not only significantly improves rendering efficiency, but also enables seamless integration with existing 3DGS acceleration kernels, further enhancing overall performance.
@@
  • A Detailed runtime analysis of the rendering pipeline showing that alpha computation in conditional alpha-blending is the performance bottleneck in 3DGS inference rendering.
  • The first use of Tensor Cores to accelerate alpha-blending, expanding the Tensor Core applicability beyond GEMM-based computations.
@@
  • A global-to-local coordinate transformation (G2L, Section 4.3) that reduces the absolute magnitudes of quadratic coordinate terms, which are sensitive to Tensor Core half-precision computation, enabling lossless rendering quality while maintaining full acceleration benefits.
2. Related Works
2.1. Fast 3D Gaussian Splatting
3DGS (Kerbl et al., 2023) has emerged as a leading real-time rendering technique with photo-realistic visual quality, finding wide applications diverse domains including autonomous driving (Zhou et al., 2024), robotics (Zhu et al., 2024), and virtual reality (Jiang et al., 2024). Current acceleration research explores two primary optimization strategies to enhance 3DGS performance.
The first strategy modifies the algorithm itself to achieve better efficiency (Ye et al., 2025; Mallick et al., 2024; Lee et al., 2024). CompactGS (Lee et al., 2024), Mini-splatting (Fang and Wang, 2024), PUP 3D-GS (Hanson et al., 2025), and gsplat (Ye et al., 2025) introduce more efficient radiance field representations and adaptive resolution primitive pruning algorithms, while Taming-3DGS (Mallick et al., 2024) presents mathematically equivalent but computationally efficient solutions for gradient computation and attribute updates, substantially accelerating training speed. Stochastic rasterization for sorting-free 3DGS (Kheradmand et al., 2025) is an another approach of improving the algorithm itself.
@@ 2.1. Fast 3D Gaussian Splatting
The second category focuses on improving computational efficiency of the original algorithm through intelligent task scheduling and CUDA kernel optimization (Feng et al., 2024; Hanson et al., 2024; Wang et al., 2024; Ye et al., 2025; Gui et al., 2024; Durvasula et al., 2023). Balanced-3DGS (Gui et al., 2024) optimizes the forward computation of rendering CUDA kernels within warps while minimizing thread idle time and maximizing resource utilization by evenly distributing tasks across computational blocks. Several works, including AdRGaussian (Wang et al., 2024), FlashGS (Feng et al., 2024), and SpeedySplat (Hanson et al., 2024), aim to reduce the number of pixels processed per Gaussian by designing more precise Gaussian-tile bounding boxes, with FlashGS (Feng et al., 2024) and SpeedySplat (Hanson et al., 2024) further computing exact Gaussian-tile intersections. DISTWAR (Durvasula et al., 2023) enhances processing efficiency through sophisticated thread management strategies, fully leveraging warp-level reductions in SM sub-cores and intelligently distributing atomic computations between SM and L2 atomic units based on memory access patterns. While these methods significantly accelerate training and rendering speed, tight coupling between algorithmic modules hinders transferability and reusability. Moreover, they fail to leverage modern hardware acceleration features such as Tensor Cores, further limiting the practical deployment and scalability of 3DGS solutions.
2.2. Tensor Cores for Non-GEMM Based Computations
[…]
@@ 2.2. Tensor Cores for Non-GEMM Based Computations
Despite this theoretical advantage, it is not a trival task to exploit Tensor Cores in arbitrary applications since Tensor Cores only support the matrix–multiply–accumulate (MMA) instruction rather than the full range of CUDA instructions set. This restriction requires algorithms to be reformulated in terms of small matrix operations, but the reward is exceptionally fast execution. Consequently, the strict requirements yet remarkable potential speed of Tensor Core operations have attracted researchers to explore adapting non-machine learning algorithms to GPU Tensor Cores, leveraging their performance potential. Previous works have successfully accelerated primitives such as reduction, scan and breadth-first search (BFS) using Tensor Cores (Dakkak et al., 2019a, b; Niu and Casas, 2025). Recently, ConvStencil (Liu et al., 2022) has managed to adapt stencil computation to TCUs, while other works have explored TCU-based implementations in applications such as HighQR (Leng et al., 2024) and TCUDB (Hu et al., 2022). However, to our best knowledge, no prior work has explored how to exploit Tensor Cores directly within a 3D Gaussian Splatting rendering pipeline.
@@
Table 1. Tensor Core & CUDA Core Comparision among various GPUs. * stands for using new sparse feature
@@
H100 (PCIe) 96 TFLOPS 800 TFLOPS — 1600 TFLOPS*
A100 78 TFLOPS 312 TFLOPS — 624 TFLOPS*
@@
RTX 4090 82.6 TFLOPS 330.3 TFLOPS — 660.6 TFLOPS*
3. Preliminaries and Observations
3.1. 3DGS Rendering Pipeline
In this paper, we focus on the official implementation of 3DGS, which is the mainstream pipeline widely used in most state-of-the-art acceleration methods.
A 3DGS model consists of a set of 3D Gaussians primitives $\mathcal{G}=\{G_{1},G_{2},\cdots,G_{P}\}$. Each Gaussian $G:\mathbb{R}^{3}\to\mathbb{R}$ is parameterized by its mean $\mu\in\mathbb{R}^{3}$, semi-definite covariance $\Sigma\in\mathbb{R}^{3\times 3}$, opacity $o\in(0,1]$, and color $c\in[0,1]^{3}$ and maps any 3D position $x\in\mathbb{R}^{3}$ to its density:
(1) $G(x)=o\mathrm{e}^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)}.$
@@ 3. Preliminaries and Observations
The preprocess stage projects 3D Gaussian $G$ into 2D planes in parallel according to the viewing transform $W$ of a given camera:
(2) $\mu^{\prime}=W\mu\in\mathbb{R}^{2},\quad\Sigma^{\prime}=JW\Sigma W^{T}J^{T}\in\mathbb{R}^{2\times 2},$
where $J=\frac{\partial{\mu^{\prime}}}{\partial\mu}$ is the Jacobian of projection $W$. $\mu^{\prime}=(\mu_{x}^{\prime},\mu_{y}^{\prime})$ and $\Sigma^{\prime}$ is the mean and the covariance of projected Gaussian $G^{\prime}$. The complexity of the preprocessing stage is $\mathcal{E}_{p}=O(P)$, where $P$ is the number of Gaussians.
@@
The screen space $\mathcal{S}$ are divided into disjoint tiles $\mathcal{T}$ with equal sizes. For each projected Gaussian, this stage identifies all the tiles it covers. Then, for each tile $\mathcal{T}\in\mathcal{S}$, the overlapped Gaussians are sorted by depth. A tile $\mathcal{T}$ coupled with a covering Gaussian $G^{\prime}$ is denoted as a splat $(\mathcal{T},G^{\prime})$. The total number of splats given by
(3) $N=\sum_{j=1}^{P}\sum_{\mathcal{T}\in\mathcal{S}}(\mathcal{T}\cap G^{\prime}_{j}\neq\emptyset)=\sum_{\mathcal{T}\in\mathcal{S}}n(\mathcal{T}).$
where $\mathbf{1}$ is the indicator function that checks whether the Gaussian covers the tile. The process of determining coverage depends on the specific algorithm used. The complexity of the sorting stage is $\mathcal{E}_{s}=O(sN)$, where $s=64$ is the coefficient in radix-sort.
Input: Tile $\mathcal{T}$, Sorted Projected Gaussians $\mathcal{G^{\prime}(\mathcal{T})}=(G^{\prime}_{1},\cdots,G^{\prime}_{n(\mathcal{T})})$
@@
foreach pixel $p\in\mathcal{T}$ in parallel do
@@
This stage renders the final color of each pixel in a tile $\mathcal{T}=(p_{1},p_{2},\cdots,p_{|\mathcal{T}|})$ in parallel as Algorithm 1. Initially, for each projected overlapping Gaussian, the renderer will perform alpha computation to get the local opacity $\alpha$ at each pixel $p=(p_{x},p_{y})\in\mathbb{R}^{2}$:
(4) $\alpha_{i,j}=o_{j}\exp{\left(-\frac{1}{2}(\mu^{\prime}_{j}-p_{i})^{T}(\Sigma^{\prime}_{j})^{-1}(\mu^{\prime}_{j}-p_{i})\right)}.$
@@
(5) $\alpha_{i,j}<\dfrac{1}{255},$
The culling process is only a conditional judgment.
Otherwise, the surviving Gaussians will perform blending and contribute the pixel’s color:
(6) $C_{i}=\sum_{j=1}^{n(\mathcal{T})}c_{j}\alpha_{i,j}T_{i,j};\quad T_{i,j}=\prod_{k=1}^{j-1}(1-\alpha_{i,k}).$
where $T_{i,j}$ denotes the transmissivity and $n(\mathcal{T})$ is the number of splats of the tile $\mathcal{T}$. If $T_{i,j}<0.0001$, the renderer will terminate the blending process for that pixel and output the color. A pixel $p$ rendered by a Gaussian $G$ is referred to as a fragment, $(p,G)$. The complexity of the alpha-blending stage is $\mathcal{E}_{a}=O(|\mathcal{T}|N)$.